OpenStudy (anonymous):

1.solve : 3^-4x+5=1/9 3^-4x+5=3^-2 -4x+5=-2 x=7/4.. 2. e^t=1000 ln e^t=ln 1000..??? i am not sure about this. haha- 3. log_2x-log_2(x-3)=1 log_2 x/x-3=1 x/x-3=2 x=2x-6 x=6...???

OpenStudy (anonymous):

For 1. awesome job! 2. e^t=1000 You are correct for taking the natural log of each side: ln(e^t)=ln(1000)=====> tln(e)=ln(1000) =====> t=ln(1000) 3. your answer is correct, i'm not sure why you have a question mark, if you could please explain so I can be of help...

OpenStudy (anonymous):

3. i wanst sure because i rechecked but it didin't match with "1" ..

OpenStudy (anonymous):

3. log_2 x - log_2 (x-3) =1 When x=6 you have log_2 (6) - log_2 (6-3) =1 log_2 (6/3) =1 2^1 does equal 6/3=2 it does check out :-)

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

you are welcome